Mathematics 244 – Linear Algebra Definitions Quiz 5 – March 5, 2004

Your Name:

Do all your work on this sheet. You have 5 minutes.

Give the definition of the following term: The matrix of the linear mapping T : V → W with respect to the bases α = {x_{1}, …, x_{n}} for V and β = {y_{1}, …, y_{m}} for W.